OpenStudy (anonymous):

A bus is moving at 22m/s [E] for 12s. Then the bus driver slows down at 1.2m/s/s [W] until the bus stops. Determine the displacement of the bus. Given: V1: 22m/s [E], V2=0, a= -1.2m/s/s [E], t=12s

OpenStudy (anonymous):

i got an answer of 177.6m but it says the answer is 4.7x10^2m

OpenStudy (anonymous):

E=positive and W=negative

OpenStudy (loser66):



OpenStudy (anonymous):

mhm

OpenStudy (loser66):

the second part: 

OpenStudy (anonymous):

yup

OpenStudy (loser66):

\(v_0 =22m/s,~~v=0~~a=-\dfrac{1}{2}m/s^2\) by \(v^2 = v_0^2 + 2aS_2\) you can find out \(S_2\) right? add \(S_1 +S_2\) you get the answer from the book

OpenStudy (anonymous):

i got 481.6m now

OpenStudy (anonymous):

would that b right still?

OpenStudy (loser66):

why? I got 465 .66 m. So, if take 2 sig.fig. you have 4.7 x 10^2 as you stated

OpenStudy (anonymous):

oh wait nvm i forgot to add 264 lol sorry im really tired atm

OpenStudy (anonymous):

i got it thanks :))))